OpenStudy (anonymous):

For what value of k will the following pair of linear equations has no solution: 3x + y = 1 and (2k-1)x + (k-1)y = 2k + 1 I am getting k=2 as solution. Can u pls confirm if it is correct. Pls give the solution with steps.

OpenStudy (anonymous):

yep its correct

OpenStudy (lalaly):

a1/a2 =b1/b2 and not equal to c1/c2 then no solution 3/(2k-1)=1/(k-1) 3(k-1)=2k-1 3k-3=2k-1 k=2

OpenStudy (anonymous):

Thanks !! I did the same thing.... Medals for you both though lalaly gave the details...☺

OpenStudy (lalaly):

and one for u coz u got it right:D

OpenStudy (anonymous):

thanks !!!!